# TEXONO Collaboration

### v1

@@ Abstract
Figure 5: Correlations between $\sigma^{\rm SR}_{\rm\chi e}$ and $\sigma^{\rm LR}_{\rm\chi e}$, showing the allowed region in the 90% C.I. from $\rm DL$ AM data [14] and in comparison with the null AM experiments at upper limits at 90% C.L. (a) with the ${\Gamma}_{tot}^{1\chi}$ analysis, where the best-fit $\rm m_{\chi}$-range is $(60{\mbox{-}}100)~{}{\rm GeV}$, and with the ${\Gamma}_{tot}^{2\chi}$ analysis, at (b) $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm GeV}$, and (c) $\rm m_{\chi}{=}(1{\pm}0.1)~{}{\rm GeV}$.
@@
On the contrary, no AM signatures were observed in the COSINE, ANAIS, XMASS, CDEX experiments. Their data are consistent with the null hypothesis. Conservative and inclusive limits are derived with single-channel (${\Gamma}_{\rm N}^{0}$) analysis -- $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ are set to be zero in Eq. Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels while the best-fit values of $\sigma^{\rm SI}_{\chi{\rm N}}$ are evaluated, allowing negative values. Limits are derived with the unified approach [29]. Same procedures apply for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ by assuming two other $\sigma$’s are zero.
@@
There are two allowed regions in ${\Gamma}_{\rm N}^{0}$ at 90% C.I. at low and high $\rm m_{\chi}$ corresponding to Na and I-recoils, respectively. With $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ are added in ${\Gamma}_{tot}^{1\chi}$, only the high $\rm m_{\chi}$ region due to I-recoils is allowed at the same significance. The best-fit solution of $\rm m_{\chi}$ is shifted from $(54.25{\pm}4.25)~{}{\rm GeV}$ in ${\Gamma}_{\rm N}^{0}$ to $(83.3{\pm}25.65)~{}{\rm GeV}$ in ${\Gamma}_{tot}^{1\chi}$. The $\rm\chi e$-channels dominate the near-threshold behavior, while the high energy (${>}3~{}\rm keV_{ee}$) spectra are defined by $\chi{\rm N}$. The two $\rm\chi e$ channels are highly correlated and share the strength of the AM counts at low energy. The limiting cases of the best-fit values in ${\Gamma}_{tot}^{1\chi}$ of $\sigma^{\rm LR}_{\rm\chi e}$ at $\sigma^{\rm SR}_{\rm\chi e}{=}0$ and $\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$ correspond to the red circles in Figures 4b&c, respectively. The exclusion contours of the null AM experiments are superimposed. The weaker bounds in $\sigma^{\rm LR}_{\rm\chi e}$ in Figure 4b originates from its sharper rise of the AM spectra due to an additional $(1/q^{2})$ dependence, so that only the low threshold data from CDEX places strong constraints. Under ${\Gamma}_{tot}^{1\chi}$, the DL AM allowed regions in Figures 4a,b&c are inter-dependent -- $\rm m_{\chi}$ is constrained to $(60{\mbox{-}}100)~{}{\rm GeV}$ by the $\rm\chi N^{SI}$ channel. The correlations of $(\sigma^{\rm LR}_{\rm\chi e},\sigma^{\rm SR}_{\rm\chi e})$ in this $\rm m_{\chi}$ range are depicted in Figure 5a, showing the DL allowed region in ${\Gamma}_{tot}^{1\chi}$ is excluded by the null AM experiments, and in particular with large margins by CDEX and XMASS combined.
@@
In the scenario of ${\Gamma}_{tot}^{2\chi}$ where $\chi{\rm N}$ and $\rm\chi e$ interactions are due to two different $\chi$’s, the constraints due to the high and low energy AM spectral components are independent. The $\rm DL$ $\rm\chi N^{SI}$ allowed regions are defined by the high-energy component and remain those of Figure 4a, which are well-excluded by the other null experiments. The AM low-energy component is dominated by the $\rm\chi e$ channels of a different $\chi$ unconstrained in $\rm m_{\chi}$. This gives rise to allowed regions in ${\Gamma}_{tot}^{2\chi}$ represented by the blue bands in Figures 4b&c for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$, respectively. As illustration, the $\rm DL$ 90% allowed C.I. for $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm and}~{}(1{\pm}0.1)~{}{\rm GeV}$ are depicted in Figure 5b&c, respectively, superimposed with the exclusion contours of the null AM data. It can be concluded that in the ${\Gamma}_{tot}^{2\chi}$ analysis, the DL AM allowed regions associated with $\rm\chi e$ are probed and rejected by the various null AM experiments. The CDEX [7] data placed more severe constraints on $\sigma^{\rm LR}_{\rm\chi e}$ where the $(1/q^{2})$ dependence favors low threshold experiments. The XMASS [6] experiment is more sensitive to $\sigma^{\rm SR}_{\rm\chi e}$ due to its large exposure. It can be projected from the results that the DL allowed regions in generic two-WIMP models are ruled out by the null AM experiments. Very little room is left to account for the DL AM data with WIMP-induced effects.

### v2

@@ III Results and Interpretations
Figure 6: Correlations between $\sigma^{\rm SR}_{\rm\chi e}$ and $\sigma^{\rm LR}_{\rm\chi e}$, showing the allowed region in the 90% C.I. from $\rm DL$ AM data [13] and in comparison with the null AM experiments at 90% C.L. upper limits. (a) with the ${\Gamma}_{tot}^{1\chi}$ analysis, where the best-fit $\rm m_{\chi}$-range is $(60{\mbox{-}}100)~{}{\rm GeV}$, and with the ${\Gamma}_{tot}^{2\chi}$ analysis, at (b) $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm GeV}$, and (c) $\rm m_{\chi}{=}(1{\pm}0.1)~{}{\rm GeV}$.
On the contrary, no AM signatures were observed in the COSINE, ANAIS, XMASS, CDEX experiments. Their data are consistent with the null hypothesis. As illustrations, the predicted AM spectra in these experiments due to the best-fit values of ($\rm\chi N^{SI}{,}\rm\chi e^{LR}{,}\rm\chi e^{SR}$) derived from DL data are depicted in Figures 4a&b. The incompatibility of the DL best-fit values with the null AM experiments can be readily visualized.
@@
Conservative and inclusive limits are derived with single-channel (${\Gamma}_{\rm N}^{0}$) analysis -- $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ are set to be zero in Eq. II.3 while the best-fit values of $\sigma^{\rm SI}_{\chi{\rm N}}$ are evaluated, allowing negative values. Limits are derived with the unified approach [29]. Same procedures apply for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ by assuming two other $\sigma$’s are zero.
III.1 Case of ${\Gamma}_{tot}^{1\chi}$
@@
There are two allowed regions in ${\Gamma}_{\rm N}^{0}$ at 90% C.I. at low and high $\rm m_{\chi}$ corresponding to Na and I-recoils, respectively. With $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ are added in ${\Gamma}_{tot}^{1\chi}$, only the high $\rm m_{\chi}$ region due to I-recoils is allowed at the same significance. The best-fit solution of $\rm m_{\chi}$ is shifted from $(54.25{\pm}4.25)~{}{\rm GeV}$ in ${\Gamma}_{\rm N}^{0}$ to $(83.3{\pm}25.65)~{}{\rm GeV}$ in ${\Gamma}_{tot}^{1\chi}$.
@@
The $\rm\chi e$-channels dominate the near-threshold behavior, while the high energy (${>}3~{}\rm keV_{ee}$) spectra are defined by $\chi{\rm N}$. The two $\rm\chi e$ channels are highly correlated and share the strength of the AM counts at low energy. The limiting cases of the best-fit values in ${\Gamma}_{tot}^{1\chi}$ of $\sigma^{\rm LR}_{\rm\chi e}$ at $\sigma^{\rm SR}_{\rm\chi e}{=}0$ and $\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$ correspond to the light red circles in Figures 5b&c, respectively. The exclusion contours of the null AM experiments are superimposed. The weaker bounds in $\sigma^{\rm LR}_{\rm\chi e}$ in Figure 5b originates from its sharper rise of the AM spectra due to an additional $(1/q^{2})$ dependence, so that only the low threshold data from CDEX places strong constraints.
Under ${\Gamma}_{tot}^{1\chi}$, the DL AM allowed regions in Figures 5a,b&c are inter-dependent -- $\rm m_{\chi}$ is constrained to $(60{\mbox{-}}100)~{}{\rm GeV}$ by the $\rm\chi N^{SI}$ channel. Figures 5b(c) shows that the low cross-section portion of $\sigma^{\rm LR}_{\rm\chi e}$($\sigma^{\rm SR}_{\rm\chi e}$) allowed regions at this $\rm m_{\chi}$ range remains unprobed by its own interactions. However, this parameter space is correlated with the high cross-section portion of its counterparts. The correlations are depicted in Figure 6a which shows the combined constraints of $(\sigma^{\rm LR}_{\rm\chi e},\sigma^{\rm SR}_{\rm\chi e})$ from the null AM experiments can reject the entire DL allowed region in ${\Gamma}_{tot}^{1\chi}$. The exclusion margins are particularly large through the combined constraints of CDEX and XMASS.
III.2 Case of ${\Gamma}_{tot}^{2\chi}$
@@
In the scenario of ${\Gamma}_{tot}^{2\chi}$ where $\chi{\rm N}$ and $\rm\chi e$ interactions are due to two different $\chi$’s, the constraints due to the high and low energy AM spectral components are independent. The $\rm DL$ $\rm\chi N^{SI}$ allowed regions are defined by the high-energy component and remain those of Figure 5a, which are well-excluded by the other null experiments. The AM low-energy component is dominated by the $\rm\chi e$ channels of a different $\chi$ unconstrained in $\rm m_{\chi}$. This gives rise to allowed regions in ${\Gamma}_{tot}^{2\chi}$ represented by the light blue bands in Figures 5b&c for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$, respectively. The ${\Gamma}_{tot}^{2\chi}$ scenario is excluded by the null AM experiments, independently in $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$. The $\rm DL$ 90% allowed C.I. in typical low and high $\rm m_{\chi}$ ranges $\left[\rm m_{\chi}{=}(100{\pm}10)~{}{\rm and}~{}(1{\pm}0.1)~{}{\rm GeV}\right]$ are depicted in Figures 6b&c, respectively, superimposed with the exclusion contours of the null AM data. The complementary roles of the experiments in probing the parameter space can be seen.
IV Conclusion
Our ${\Gamma}_{tot}^{2\chi}$ analysis and results indicate that the DL AM allowed regions associated with $\rm\chi e$ are probed and rejected by the various null AM experiments. The CDEX [7] data placed more severe constraints on $\sigma^{\rm LR}_{\rm\chi e}$ where the $(1/q^{2})$ dependence favors low threshold experiments. The XMASS [6] experiment is more sensitive to $\sigma^{\rm SR}_{\rm\chi e}$ due to its large exposure.
It can be projected that the DL allowed regions in generic two-WIMP models interacting via the $\chi{\rm N}$ and $\rm\chi e$ channels independently are ruled out by the null AM experiments. Very little room is left to account for the DL AM data with WIMP-induced $\chi{\rm N}{+}\rm\chi e$ effects.
V Acknowledgments